# ResSVD: Residual Compensated SVD for Large Language Model Compression

### v2

@@ 4.3 Ablation Study
In this section, we present extensive ablation study results to assess the effectiveness and robustness of ResSVD from three key perspectives. ① Effectiveness of residual compensation and partial-layer compression: We evaluate the individual contributions of the residual compensation mechanism and partial-layer compression. ③ Impact of layer selection strategy: We show that restricting compression to the last few layers mitigates error accumulation and consistently improves performance. ③ Impact of calibration data: We analyze the effects of both calibration sample size and dataset choice on the performance of compressed models.
Effectiveness of residual compensation and partial-layer compression. As shown in Table 4, we conduct ablation studies to evaluate the individual contributions of residual compensation (REC) and partial-layer compression (PLC). Specifically, the left table presents the results for REC under multiple compression ratios. Incorporating REC consistently reduces perplexity across all settings, and ResSVD achieves the best performance across all datasets and baselines. The right table reports the results for PLC. Without PLC, perplexity on WikiText-2 [29] increases slightly, indicating the effect of cumulative error. In contrast, applying PLC consistently enhances performance, further amplifying the benefits of REC. These results confirm that PLC is essential for mitigating error propagation. More ablation results are provided in Appendix A.3.
Impact of layer selection strategy. We compare the performance of LLaMA-7B [38] under 20% compression ratio with different layer selection strategies. The results are shown in Table 5. We observe that the performance of the compressed model varies significantly depending on which layers are selected for compression. The worst performance is observed when compressing the first 8 layers, primarily because the compression ratio per layer becomes excessively high, and error accumulation begins from the first layer. Across all configurations, compressing the last 16 layers yields the best overall performance. However, as more layers are compressed, the performance gradually declines, which can be attributed to increased cumulative approximation error.
Impact of calibration data.   The selection of calibration data plays a critical role in post-training compression. To assess this, we vary both the number of calibration samples and the choice of calibration datasets. Figure 4 shows the perplexity on WikiText-2 [29] and average zero-shot accuracy under 20% compression ratio. While performance slightly improves with more calibration samples, the gains are modest, indicating that ResSVD remains robust even with limited calibration data. Moreover, Table 6 shows that the best results are achieved when the calibration and evaluation datasets align, suggesting that domain consistency enhances compression quality.
@@ 4.4 Efficiency Results
ResSVD not only preserves competitive model performance but also delivers significant inference speedup on hardware. We evaluate the throughput of compressed models on an NVIDIA A100 GPU, with results shown in Figure 5. Models compressed by ResSVD consistently achieve faster inference than those compressed by SVD-LLM [41]. Moreover, the speedup becomes increasingly significant as the batch size grows, indicating that ResSVD scales more efficiently under larger workloads. These findings highlight the practical efficiency of ResSVD in enabling faster inference while maintaining accuracy, making it well-suited for deployment in resource-constrained environments.
Table 4: Ablation results of residual compensation (REC) and partial-layer compression (PLC) on LLaMA-7B [38] under compression ratios 20% and 30%. More ablation results are in Table 8.
(a) REC
@@
20% SVD 20082.86 20338.96 18784.20
ASVD [44] 9.27 15.09 13.68
SVD-LLM [41] 7.89 16.54 15.92
ResSVD 7.93 13.68 13.72
ResSVD 7.47 12.27 12.22
30% SVD 13155.97 17354.46 21012.91
ASVD [44] 222.98 586.79 148.79
SVD-LLM [41] 9.52 28.97 26.38
ResSVD 9.60 26.12 20.97
ResSVD 9.52 20.31 18.29
(b) PLC
@@
20% SVD 20082.86 20338.96 18784.20
ASVD [44] 9.27 15.09 13.68
SVD-LLM [41] 7.89 16.54 15.92
ResSVD 8.27 16.00 15.78
ResSVD 7.47 12.27 12.22
30% SVD 13155.97 17354.46 21012.91
ASVD [44] 222.98 586.79 148.79
SVD-LLM [41] 9.52 28.97 26.38
ResSVD 10.11 24.49 23.82
ResSVD 9.52 20.31 18.29
Table 5: Performance of compressed LLaMA-7B [38] with various layer selection strategies.
@@
First 8 80% 16.56 37.58 46.16 0.33
Last 8 80% 10.69 19.26 16.06 0.38
Last 12 53% 8.49 14.04 13.50 0.44
Last 16 40% 7.47 12.27 12.22 0.47
Last 20 32% 8.07 14.19 13.68 0.46
Last 24 27% 8.08 14.66 13.94 0.45
Last 28 23% 8.21 16.04 14.98 0.45
Last 32 20% 8.27 16.00 15.78 0.44
Table 6: Perplexity ($\downarrow$) and zero-shot average accuracy ($\uparrow$) of LLaMA-7B [38] compressed by ResSVD under 20% compression ratio while using different calibration datasets.
@@
WikiText-2 [29] 7.47 12.27 12.22 0.47
PTB [28] 9.86 9.54 12.70 0.45
C4 [32] 10.05 12.42 10.94 0.47
Figure 4: Impact of the number of calibration data samples on LLaMA-7B [38] under 20% compression ratio. (Left) Perplexity ($\downarrow$). (Right) Average accuracy ($\uparrow$).
Figure 5: Throughput of LLaMA-7B [38] and its 40% compressed versions. The sequence length is 32.
5 Conclusion
In this work, we propose ResSVD, a novel post-training SVD-based compression method for large language models. ResSVD effectively leverages the residual matrix resulting from SVD truncation to reduce the truncation loss and enhance layer-wise reconstruction accuracy. Furthermore, it selectively compresses only the last few layers of the model under a fixed overall compression ratio, thereby significantly mitigating error propagation across the model. Extensive experiments across various LLM families and benchmark datasets demonstrate that ResSVD consistently outperforms existing SVD-based baselines across various settings. These results highlight the effectiveness and generalizability of ResSVD in enabling efficient LLM deployment.
Appendix A Appendix
A.1 Pseudocode
Algorithm 2 and Algorithm 3 present the pseudocode for partial-layer compression and residual compensation, respectively. Algorithm 2 identifies the optimal number of last layers $k^{\star}$ to compress, along with their corresponding layer compression ratio $R_{l}^{\star}$. Specifically, for a model with $N$ layers, we iterate over candidate values of $k$ using a step size $s$ and compute the associated $R_{l}$ for each. For each candidate configuration, we invoke Algorithm 3 to perform the compression, after which we compute the final-layer error relative to the original model. The configuration yielding the lowest final-layer error is selected as the optimal compression setting.
@@ A.1 Pseudocode
0:  Original LLM: $M$, number of model layers: $N$, overall compression ratio: $R_{o}$, intermediate rank: $r_{1}$, step: $s$
@@
1:  $CD\leftarrow$ Randomly collect calibration samples from the dataset
2:  $\mathrm{Set}_{\boldsymbol{W}}\leftarrow M$triangleright Extract the set of weight matrices in $M$
3:  $\mathrm{Set}_{k}\leftarrow\{\,k\mid k=s,2s,\dots,N-s\}$triangleright Obtain candidate compressed layer numbers
4:  $Lowest\_Err\leftarrow+\infty$
5:  for each $k$ in $\mathrm{Set}_{k}$ do
6:     $R_{l}\leftarrow(N\cdot R_{o})/k$triangleright Calculate corresponding layer compression ratio
7:     if $R_{l}\geq 0$ then
8:        $M_{TMP}\leftarrow$ Deep_Copy$(M)$
9:        for $i\in[N-k+1,N]$ do
10:           $W_{i}\leftarrow$ $\mathrm{Set}_{\boldsymbol{W}}(i)$, $S_{i}\leftarrow$ Whitening$(W_{i},CD)$triangleright Initialize weight matrices to compress
11:           $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow$ Residual Compensation($W_{i}S_{i}$, $k$, $R_{l}$, $r_{1}$)
@@
14:        $Err\leftarrow$ Cal_Error($M,M_{TMP},CD,Layer(N)$) $\triangleright$ Calculate the last layer output error
15:        if $Err<Lowest\_Err$ then
16:           $Lowest\_Err\leftarrow Err$triangleright Search for the lowest layer-wise error
17:           $k^{\star}\leftarrow k$
18:           $R_{l}^{\star}\leftarrow R_{l}$
@@
22:  return  $k^{\star},R_{l}^{\star}$
Algorithm 3 Pseudocode of Residual Compensation
0:  Original LLM: $M$, weight matrix: $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, number of model layers: $N$, overall compression ratio: $R_{l}$, step: $s$, intermediate rank: $r_{1}$
1:  $CD\leftarrow$ Randomly collect calibration samples from the dataset
2:  $\mathrm{Set}_{\boldsymbol{W}}\leftarrow M$triangleright Extract the sets of weight matrices in $M$ to compress
3:  $\mathrm{Set}_{\boldsymbol{S}}\leftarrow$ Whitening($\mathrm{Set}_{\boldsymbol{W}}$, $CD$), $\mathrm{Set}_{\boldsymbol{W^{\prime}}}\leftarrow\emptyset$triangleright Initialize sets of weight matrices
4:  $k^{\star}$, $R_{l}^{\star}\leftarrow$ Partial-layer Compression($M$, $N$, $R_{o}$, $s$)
5:  $r=m\cdot n\cdot(1-R_{l}^{\star})/(m+n)$, $r_{2}=r-r_{1}$
6:  for $i\in[N-k^{\star}+1,N]$ do
7:     $\boldsymbol{W}_{i}\leftarrow\mathrm{Set}_{\boldsymbol{W}}(i)$, $\boldsymbol{S}_{i}\leftarrow\mathrm{Set}_{\boldsymbol{S}}(i)$triangleright Extract whitening matrices of current weight matrices
8:     $\boldsymbol{U}_{i,r_{1}}$, $\boldsymbol{\Sigma}_{i,r_{1}}$, $\boldsymbol{V}_{i,r_{1}}^{T}$leftarrow $\boldsymbol{W}_{i,r_{1}}$leftarrow SVD_Trunc($\boldsymbol{W}_{i}\boldsymbol{S}_{i}$) $\triangleright$ Apply the first SVD truncation
9:     $\boldsymbol{R}_{i}$leftarrow Cal_Res($\boldsymbol{W}_{i}$, $\boldsymbol{W}_{i,r_{1}}$) $\triangleright$ Calculate residual matrices
10:     $\boldsymbol{U}_{i,r_{2}}$, $\boldsymbol{\Sigma}_{i,r_{2}}$, $\boldsymbol{V}_{i,r_{2}}^{T}$leftarrow $\boldsymbol{R}_{i,r_{2}}$leftarrow SVD_Trunc($\boldsymbol{R}_{i}$) $\triangleright$ Apply the second SVD truncation
11:     $\hat{\boldsymbol{U}}_{i,r_{1}}$leftarrow Mul($\boldsymbol{U}_{i,r_{1}}$, $\sqrt{\boldsymbol{\Sigma}_{i,r_{1}}}$), $\hat{\boldsymbol{V}}_{i,r_{1}}$leftarrow Mul($\sqrt{\boldsymbol{\Sigma}_{i,r_{1}}}$, $\boldsymbol{V}_{i,r_{1}}^{T}$) $\triangleright$ Absorb singular values
12:     $\hat{\boldsymbol{U}}_{i,r_{2}}$leftarrow Mul($\boldsymbol{U}_{i,r_{2}}$, $\sqrt{\boldsymbol{\Sigma}_{i,r_{2}}}$), $\hat{\boldsymbol{V}}_{i,r_{2}}$leftarrow Mul($\sqrt{\boldsymbol{\Sigma}_{i,r_{2}}}$, $\boldsymbol{V}_{i,r_{2}}^{T}$) $\triangleright$ Absorb singular values
13:     $\hat{\boldsymbol{U}}_{i,r}$leftarrow Comb($\hat{\boldsymbol{U}}_{i,r_{1}}$, $\hat{\boldsymbol{U}}_{i,r_{2}}$), $\hat{\boldsymbol{V}}_{i,r}$leftarrow Comb($\hat{\boldsymbol{V}}_{i,r_{1}}$, $\hat{\boldsymbol{V}}_{i,r_{2}}$) $\triangleright$ Combine weight matrices $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$leftarrow ($\hat{\boldsymbol{U}}_{i,r}$, $\hat{\boldsymbol{V}}_{i,r}$) $\cup$ $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$triangleright Append decomposed weight matrices
14:  end for
15:  return  $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$
A.2 Overall Performance of LLaMA-2-7B
We evaluate the perplexity and zero-shot accuracy of LLaMA-2-7B [39] to assess the effectiveness of ResSVD, with results reported in Table 7. On the three language modeling datasets, ResSVD outperforms all baselines across the evaluated compression ratios, except at the 60% compression ratio on WikiText-2 [29]. In particular, compared to the existing best baseline, the relative improvement ranges from 3% to 75%. For compression ratios of 40%, 50%, and 60%, we also include results from the recent method AdaSVD [22], using values reported in their paper due to the lack of an open-source codebase. ResSVD continues to outperform AdaSVD across these settings. Moreover, on zero-shot accuracy across seven common sense reasoning datasets, ResSVD achieves the best performance on most benchmarks, with an average improvement of up to 16%. These results further demonstrate the robustness and overall effectiveness of ResSVD.
Table 7: Overall performance of LLaMA-2-7B [39] compressed by ResSVD and baselines under 20% to 60% compression ratios ("Ratio"), including performance on three language modeling datasets (measured by perplexity ($\downarrow$)) and zero-shot performance on seven common sense reasoning datasets (measured by individual and average accuracy ($\downarrow$)). The best results are marked in bold. Blue arrows within parentheses highlight the relative improvement over the existing second-best baseline. NaN denotes evaluation failure due to numerical instability.
@@ A.2 Overall Performance of LLaMA-2-7B
- Original 5.47 26.84 7.28 0.31 0.69 0.67 0.56 0.40 0.78 0.27 0.53
ASVD [44] 9.56 120.74 12.85 0.25 0.59 0.61 0.46 0.32 0.72 0.24 0.45
SVD-LLM [41] 8.37 139.68 20.18 0.23 0.50 0.59 0.39 0.26 0.65 0.23 0.41
ResSVD 7.63 ($\downarrow$9%) 45.37 ($\downarrow$62%) 14.73 ($\downarrow$8%) 0.28 0.61 0.65 0.50 0.35 0.73 0.26 0.48 ($\uparrow$7%)
ASVD [44] 984.03 NaN NaN 0.15 0.27 0.51 0.26 0.22 0.53 0.20 0.31
SVD-LLM [41] 10.66 292.90 34.96 0.21 0.42 0.55 0.34 0.22 0.60 0.23 0.37
ResSVD 10.32 ($\downarrow$3%) 73.04 ($\downarrow$75%) 23.28 ($\downarrow$33%) 0.23 0.51 0.62 0.42 0.29 0.68 0.24 0.43 ($\uparrow$ 16%)
ASVD [44] NaN NaN NaN 0.15 0.25 0.50 0.26 0.22 0.52 0.18 0.30
SVD-LLM [41] 16.11 717.34 61.96 0.16 0.35 0.55 0.30 0.20 0.57 0.23 0.34
AdaSVD [22] 14.76 304.62 56.98 0.19 0.41 0.58 0.32 0.23 0.58 - -
ResSVD 14.17 ($\downarrow$4%) 136.32 ($\downarrow$55%) 43.19 ($\downarrow$24%) 0.20 0.43 0.57 0.35 0.24 0.63 0.23 0.38 ($\uparrow$12%)
ASVD [44] NaN NaN NaN 0.13 0.26 0.50 0.25 0.23 0.50 0.20 0.30
SVD-LLM [41] 27.19 1775.52 129.71 0.14 0.30 0.50 0.28 0.20 0.54 0.23 0.31
AdaSVD [22] 25.58 593.14 113.84 0.15 0.34 0.54 0.29 0.20 0.56 - -
ResSVD 24.26 ($\downarrow$5%) 286.24 ($\downarrow$52%) 100.34 ($\downarrow$12%) 0.14 0.35 0.55 0.31 0.22 0.59 0.22 0.34 ($\uparrow$10%)
ASVD [44] NaN 19581.17 NaN 0.15 0.25 0.50 0.25 0.23 0.52 0.12 0.29
SVD-LLM [41] 54.19 3442.74 263.02 0.14 0.26 0.50 0.27 0.20 0.53 0.21 0.30
AdaSVD [22] 60.08 2137.28 294.26 0.12 0.27 0.50 0.27 0.20 0.53 - -
ResSVD 58.88 (-) 991.48 ($\downarrow$54%) 255.70 ($\downarrow$3%) 0.13 0.29 0.52 0.28 0.21 0.55 0.22 0.31 ($\uparrow$3%)
A.3 More Ablation Results
The ablation results for LLaMA-7B [38] under 40%, 50%, and 60% compression ratios are presented in Table 8. A similar trend to that observed at 20% and 30% compression holds here: incorporating both REC and PLC leads to a substantial reduction in perplexity across all settings. Moreover, ResSVD continues to outperform all existing baselines, further validating the effectiveness of these two components under higher compression ratios.
Table 8: Ablation results of Residual Compensation (REC) and Partial-layer Compression (PLC) on LLaMA-7B [38] under compression ratios 40%, 50% and 60%.
(a) REC

### v3

@@ 4.3 Ablation Study
We present several ablation studies to assess the robustness of ERC-SVD. ① Effectiveness of residual compensation (REC) and partial-layer compression (PLC): We assess the individual contributions of REC and PLC. ② Impact of residual compensation factor: We conduct experiments to examine how the choice of $\beta$ influences the performance. ③ Impact of calibration data: We analyze the effects of calibration dataset selection on compressed model performance in Appendix A.3.4.
Effectiveness of residual compensation and partial-layer compression.
We present results on C4 [raffel2020exploring] and seven zero-shot reasoning and understanding tasks, as shown in Table 4. It can be observed that either applying REC or PLC alone has already improved performance. Combining them leads to further improvements, making the performance gap even more significant. Results of other compression ratios are presented in Table 8.
Impact of residual compensation factor.
The residual compensation factor $\beta$ is a hyperparameter in this work, we set $\beta=0.05$ across all experiments. To examine its impact on model performance, we compress LLaMA-2-7B under 30% compression ratio while varying the value of $\beta$. The results in Table 5 show that the compressed models achieve comparable performance across different $\beta$ values, demonstrating the robustness of our method with respect to this hyperparameter.
4.4 Compatibility with Quantization
SVD-based LLM compression methods and quantization are two orthogonal techniques. To demonstrate that our method can be integrated with quantization, we adopt GPTQ [frantar2022gptq] to quantize LLaMA-2-7B compressed by our method and SVD-LLM. As shown in Table 12, our method integrates seamlessly with quantization, achieving superior performance compared to SVD-LLM.
@@ 4.5 Efficiency Results
ERC-SVD not only preserves competitive model performance but also achieves substantial inference speedup on hardware. We evaluate the throughput of compressed models on an NVIDIA A100 GPU, shown in Figure 5. Models compressed by ERC-SVD consistently deliver faster inference than the original model. The speedup grows more pronounced as the batch size grows, indicating that ERC-SVD scales more efficiently under larger workloads. Additionally, Appendix A.6 provides further analysis on how ERC-SVD reduces computational complexity.
5 Conclusion
In this work, we propose ERC-SVD, a novel post-training SVD-based compression method for LLMs, formulated from an error-controlled perspective. ERC-SVD effectively leverages the residual matrix resulting from SVD truncation to reduce the truncation loss and enhance layer-wise reconstruction accuracy. Furthermore, it selectively compresses only the last few layers of the model, thereby significantly mitigating error propagation. Extensive experiments across various LLM families and benchmark datasets demonstrate that ERC-SVD consistently outperforms existing SVD-based baselines under various settings. These results highlight the effectiveness and generalizability of ERC-SVD in enabling efficient LLM deployment.
Appendix A Appendix
Here, we provide further details that are not discussed in the main paper and include extra experimental results. The appendix is structured as follows:
A.1 Pseudocode
Algorithm 2 and Algorithm 3 present the pseudocode for residual compensation and partial-layer compression, respectively. Algorithm 3 identifies the optimal number of last layers $k$ to compress, along with their corresponding layer compression ratio $R_{l}$. Specifically, for a model with $N$ layers, we iterate over candidate values of $k^{\prime}$ using a step size $s$ and compute the associated $R_{l}^{\prime}$ for each. For each candidate configuration, we invoke Algorithm 2 to perform the compression, after which we compute the final-layer error relative to the original model. The configuration yielding the lowest final-layer error is selected as the optimal compression setting.
Algorithm 2 Pseudocode of Residual Compensation
0: Original LLM: $M$, weight matrix: $\boldsymbol{W}_{i}\in\mathbb{R}^{m\times n}$, layer compression ratio: $R_{l}$, residual compensation rank: $r_{r}$
@@ A.1 Pseudocode
[…]
12:return $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)$
@@
0: Original LLM: $M$, number of model layers: NN, overall compression ratio: $R_{o}$, step: $s$
@@
[…]
5:for each $k^{\prime}$ in $\mathrm{Set}_{k^{\prime}}$ do
6:  $R_{l}^{\prime}\leftarrow(N\cdot R_{o})/k^{\prime}$ $\triangleright$ Calculate corresponding layer compression ratio
7:  if $R_{l}^{\prime}\geq 0$ then
8:   $M_{TMP}\leftarrow$ Deep_Copy(M)(M)
9:   for $i\in[N-k^{\prime}+1,N]$ do
10:    $\textbf{W}_{i}\leftarrow$ $\mathrm{Set}_{\boldsymbol{W}}(i)$ $\triangleright$ Initialize weight matrices to compress
11:    $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow$ Residual Compensation($M$, $\textbf{W}_{i}$, $R_{l}^{\prime}$, $r_{r}$)
@@
[…]
15:   if Err<Lowest_ErrErr<Lowest\_Err then
16:    $Lowest\_Err\leftarrow Err$ $\triangleright$ Search for the lowest layer-wise error
17:    $k\leftarrow k^{\prime}$
18:    $R_{l}\leftarrow R_{l}^{\prime}$
@@
22:return $k,R_{l}$
A.2 Layer-wise Error Comparison
As discussed in Section 3.2, the same observation also holds here. We conduct experiments on Mistral-7B [jiang2023mistral] and Vicuna-7B [chiang2023vicuna], and the results are presented in Figure 4.
Figure 4: Layer-wise error comparison between the original model, Mistral-7B, and Vicuna-7B compressed by ERC-SVD with different layer selection strategies on WikiText-2. The overall compression ratio is 20%, and all layer selection strategies strictly adhere to the compression constraint.
A.3 Additional Results
In this section, we present additional results from four perspectives: Section A.3.1 reports zero-shot accuracy on larger-scale LLMs, Section A.3.2 reports zero-shot accuracy across multiple LLM families, Section A.3.3 presents ablation results under compression ratios 40%, 50%, and 60%, and Section A.3.4 demonstrates the impact of calibration data.
A.3.1 Results on Larger-scale LLMs
In Table 6, we present zero-shot accuracy results for LLaMA-13B [touvron2023llama], LLaMA-30B [touvron2023llama], LLaMA-2-13B [touvron2023llama2], OPT-13B [zhang2022opt], and OPT-30B [zhang2022opt]. Across these evaluation datasets, ERC-SVD consistently outperforms SVD-LLM, with only one minor exception.
Table 6: Zero-shot accuracy results across larger-scale LLMs under 20% compression ratio.
@@ A.3.1 Results on Larger-scale LLMs
LLaMA-13B SVD-LLM 0.27 0.62 0.66 0.46 0.33 0.72 0.26 0.47
ERC-SVD 0.28 0.69 0.69 0.54 0.39 0.75 0.26 0.52
LLaMA-30B SVD-LLM 0.29 0.72 0.73 0.51 0.38 0.74 0.27 0.51
ERC-SVD 0.30 0.73 0.73 0.59 0.43 0.78 0.29 0.55
LLaMA-2-13B SVD-LLM 0.27 0.63 0.66 0.44 0.31 0.71 0.26 0.47
ERC-SVD 0.31 0.67 0.68 0.47 0.38 0.74 0.27 0.50
OPT-13B SVD-LLM 0.24 0.62 0.64 0.47 0.30 0.72 0.24 0.46
ERC-SVD 0.24 0.63 0.65 0.45 0.30 0.73 0.24 0.46
OPT-30B SVD-LLM 0.29 0.67 0.65 0.51 0.33 0.75 0.25 0.49
ERC-SVD 0.28 0.68 0.67 0.50 0.34 0.76 0.25 0.49
A.3.2 Results on Multiple LLM Families
We also evaluate the zero-shot accuracy of OPT-6.7B [zhang2022opt], Mistral-7B [jiang2023mistral], and Vicuna-7B [chiang2023vicuna]. Results are shown in Table 9, and our method consistently outperforms SVD-LLM across these diverse LLM architectures. In addition, we conduct experiments on Qwen-3-8B [yang2025qwen3], with perplexity and zero-shot accuracy results reported in Table 7.
Table 7: Performance of Qwen-3-8B under 20% to 60% compression ratios (“Ratio”).